FOUR DIFFERENT CRIME VICTIMS LISTENED TO VOICE RECORDINGS OF SEVEN DIFFERENT MEN. IF THE VOICE IDENTIFICATIONS WERE MADE BY RANDOM GUESS, FIND THE PROBABILITY THAT ALL VICTIMS WOULD SELECT A DIFFERENT VOICE RECORDING
Answer #1

Out of 7 different voices, the probability that a victim selects a different voice recording(not identifying the actual person) =6/7

So, the probability that all 4 victims select a different voice recording =(6/7)*(6/7)*(6/7)*(6/7) =(6/7)4 =0.5398